OpenStudy (anonymous):

let, I=∫tan^-1x dx =tan^-1x ∫1 dx-∫{d/dx tan^-1x ∫1 dx} dx =x tan^-1x -∫x/1+x^2 dx =x tan^-1x- ????

OpenStudy (anonymous):

@amistre64 help

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@dumbcow :(

OpenStudy (dumbcow):

ok trying to follow your work lol looks like you did integration by parts and you are stuck at integrating x/(1+x^2) use substitution: w = 1 + x^2 dw = 2x dx ---> dx = dw/2x "the x's will cancel"

OpenStudy (anonymous):

so x tan^-1x -∫dz/2/z =x tan^-1x-1/2∫1/z dz =x tan^-1x-1/2 ln z =x tan^-1x-1/2 ln (1+x^2) ans???

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

help bro :'(

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

yep :) use this to check answers http://www.wolframalpha.com/input/?i=integrate+arctan%28x%29+dx

OpenStudy (anonymous):

but if i use the method of substitution in integration by parts then what if the teacher is gonna cut my number :(

OpenStudy (anonymous):

thank you i will definitely take help from that :) thank you very very much :) bro

OpenStudy (dumbcow):

the teacher shouldn't....many integrals require multiple methods to finish the substitution is done within the integration by parts so i think you are ok

OpenStudy (anonymous):

thank you :)